# Tunable topological phase in 2D ScV6Sn6 kagome material

### v1

@@ III RESULTS AND DISCUSSION
Figure 3: (a) Topologically protected edge state in the TSM phase. The white dashed lines in the inset highlight the closing of the topological gap in the surface at ‘w‘. (b) Berry curvature distributions of FM 2D ScV6Sn6 in the first Brillouin zone, with red regions marking Berry curvature singularities. (c)-(d) Fermi surface of the 110 slab: (c) without external fields and (d) with an electric field of 0.4 eV/Å, highlighting changes in Fermi pockets at the critical E-field. (e) The topological phase transition from the TSM phase ($|C|=1$) to the trivial phase ($|C|=0$), with the white dashed line showing the transition regime. (f) Edge states of FM 2D ScV6Sn6 (110) at critical E-field of 0.4 eV/Å, where red lines enclose bands containing nodal points, and the dotted line indicates the Fermi level.
In Figure 3, we investigate the influence of external perpendicular electric fields on the topological properties of 2D ScV6Sn6. The edge states (Figures 3(a)-(b)) corresponding to the Weyl point at the Fermi level, along with the Berry curvature distribution at zero field, are characterized by a band crossing within the topological band between the $M$ and $K$ high-symmetry points in the Brillouin zone. Additionally, the Berry curvature hotspots (Figure 3(b)) correspond to Weyl point pairs located at the Fermi level and slightly below it in the valence band. Upon applying external electric fields, distinct modifications in the electronic band structure emerge, leading to a transition from a topologically nontrivial to a trivial phase. Under small perturbative fields, however, the overall topology of the bands remains largely intact. Variations in the chirality of the Weyl points are observed at different field strengths, but the system retains its topological phase until a critical field strength of $\sim$0.4 eV/Å. At this critical electric field, the edge states begin to open up, corresponding to a spin-orbit coupling-induced bulk band gap. As shown in Figure 3(f), the band crossings that were present at zero field either open or vanish entirely, confirming a transition to a topologically trivial state. This is further supported by the disappearance of the topologically protected edge states and a corresponding switch in the Chern number from 1 to 0 (Figure 3(e)), indicating the system’s transition to a trivial metallic phase. Beyond this critical field, stronger electronic coupling and coherence become evident, as shown by the fully interconnected Fermi pockets (Figure 3(d)). These interconnected pockets suggest coherent charge carrier motion in the trivial metallic state after the topological transition. A similar electric-field-induced topological transition has been observed in MnBi2Te4 monolayers and other 2D materials, such as MoS2 and WSe2, where external fields modulate their topological states [59, 24, 60].
@@
In our simulations, we observed fluctuations in the Chern number at smaller electric fields, likely caused by modifications in the Berry curvature, driven by electric-field-induced alterations in the band structure. Although the precise mechanism behind these fluctuations remains unclear, similar behavior has been reported in MnBi2Te4, where the Chern number can be toggled by applying an electric field, with transitions between discrete values such as 1 and 2 at field strengths of around 1 to 2 eV/Å [61]. Additionally, thin films near the topological critical point have exhibited fluctuations in the Chern number between 0, 1, and -1 as the electric field changes, eventually stabilizing at zero beyond approximately 2.5 eV/Å [62]. Rasouli et al. also observed similar electric-field-induced changes in the Chern number for 2D MnO2, with stabilization at zero occurring beyond a certain threshold.
@@
In contrast, the application of small magnetic fields (B-fields) does not alter the topology of the bands. The Chern number remains unchanged regardless of the direction or magnitude of the applied magnetic field. This stability is due to the robustness of the topologically protected edge states, preserved by the system’s geometry and time-reversal symmetry (TRS). As moderate magnetic fields are insufficient to break this symmetry, the topological phase remains protected, similar to the role of spin-orbit coupling (SOC), which is also unable to break the symmetry [63]. Moreover, the weak coupling between the magnetic field and the electronic states in this material results in minimal effects on the band structure and topological properties. The stability of the Fermi surface under both in-plane and out-of-plane magnetic fields up to 2.1 T (not shown) further supports this robustness. The Fermi surface retains its 2D character, characterized by large hexagonal V-$d$ states and smaller circular Sc-$d$ states, as illustrated in Figures 3(b)-(c). Importantly, the shape of the Fermi surface remains consistent across different magnetic field strengths, indicating that the electronic structure is preserved under external magnetic fields. These findings are consistent with studies on other magnetic topological materials. For example, in MnBi2Te4, although magnetic fields induce transitions between antiferromagnetic (AFM) and ferromagnetic (FM) states, the Fermi surface remains largely unaffected, and no topological transition is observed [64]. Similarly, in Cr-doped Bi2Se3 and 2D materials such as CrI3 and Cr2Ge2Te2, minimal topological changes occur under magnetic fields despite notable effects on magnetic properties [65, 66].
@@ IV Conclusions
We have investigated the 2D kagome metal ScV6Sn6 and demonstrated that it exhibits topologically protected edge states and a Chern number of $|\textit{C}|=1$ at the Fermi energy, confirming the presence of a Weyl semimetallic phase. This positions ScV6Sn6 among the few vanadium-based kagome compounds with nontrivial topology and broken time-reversal symmetry. Our analysis of edge states originating from gapped Weyl points near the Fermi energy reveals that the material retains metallic behavior at the surface despite being gapped in the bulk, confirming its topological nature. We further show that while magnetic fields do not affect its topology, an electric field of 0.4 eV/Å induces a transition from a topological semimetal to a trivial metallic phase, with the edge states vanishing. This electrically-induced transition, consistent with observations in other 2D materials, underscores the tunability of ScV6Sn6’s topological properties. Additionally, our calculations reveal a large anomalous Hall effect of $\sim$257 $\Omega^{-1}$cm-1, driven by intrinsic Berry curvature and strong spin-orbit coupling, aligning with other 2D ferromagnetic Dirac-like materials. These findings highlight ScV6Sn6 as a pivotal material in 2D kagome systems, with strong potential for modulating topological properties in $d$-orbital-dominated structures. Its robust ferromagnetic ordering, tunable topological phases, and spin-orbit coupling make it an ideal candidate for advanced electronic and spintronic applications. The ability to control topological transitions through electric fields not only enhances our understanding of topological quantum materials but also paves the way for innovations in quantum computing and material design.
@@ V Acknowledgments
This research is supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences under Award DOE-SC0024099 (algorithmic development and first-principles calculations) and the Lehigh Core grant (initial high-throughput framework development). SKD was supported by the start up grant of Bitan Roy from Lehigh University. Computational resources were provided by Lehigh University Research Computing Infrastructure. We also acknowledge useful discussions with Benjamin Wieder.

### v2

@@ III RESULTS AND DISCUSSION
Figure 3: (a) Topologically protected edge state in the TSM phase. The white dashed lines in the inset highlight the closing of the topological gap in the surface at ‘w‘. (b) Berry curvature distributions of FM 2D ScV6Sn6 in the first Brillouin zone, with red regions marking Berry curvature singularities. (c)-(d) Fermi surface of the 110 slab: (c) without external fields and (d) with an electric field of 0.4 eV/Å, highlighting changes in Fermi pockets at the critical E-field. (e) The topological phase transition from the TSM phase ($|C|=1$) to the trivial phase ($|C|=0$), with the white dashed line showing the transition regime. (f) Edge states of FM 2D ScV6Sn6 (110) at a critical E-field of 0.4 eV/Å, where red lines enclose bands containing nodal points. The dashed green box represents the position of the Weyl point $w$, while the dotted black line indicates the Fermi level.
In ferromagnetic materials, the total anomalous Hall effect (AHE) conductivity, $\sigma_{H}^{A}$, arises from three main contributions: the intrinsic contribution $\sigma_{int}$, skew scattering $\sigma_{sc}$, and the side-jump mechanism $\sigma_{sj}$ chen2021large ; nagaosa2010anomalous . Mathematically, $\sigma_{H}^{A}$ is expressed as $\sigma_{H}^{A}=\sigma_{int}+\sigma_{sc}+\sigma_{sj}$. The intrinsic contribution, $\sigma_{int}$, depends on the material’s band structure and is independent of scattering rates, while $\sigma_{sc}$ and $\sigma_{sj}$ originate from extrinsic scattering processes caused by impurities or disorder. We calculated the energy-dependent intrinsic AHC ($\sigma_{int}$) using the Berry curvature derived from band calculations. Figure 2(d) shows the AHC plot, with a peak near the Fermi energy, reaching 257 $\Omega^{-1}\text{cm}^{-1}$. The strong SOC and ferromagnetic ordering in 2D ScV6Sn6 enhance the Berry curvature, which acts as an effective magnetic field in momentum space, contributing to the large AHC. The peak at $\sim$0.33 eV near the Fermi energy can be attributed to the proximity of band crossings to the Fermi energy. The large anomalous Hall effect is believed to arise from the Berry curvature hotspots near the Fermi level. SOC-induced small gaps act as these Berry curvature hotspots, which in turn lead to the observed AHC. Currently, there are limited reports on AHC in 2D materials, making our findings particularly significant. This study provides valuable insights into the anomalous Hall effect in 2D kagome materials and contributes to the broader understanding of topological phenomena in these systems.
Previous studies on 2D ferromagnetic Dirac materials have reported intrinsic anomalous Hall conductivity (AHC) values reaching up to approximately 300 $\Omega^{-1}\text{cm}^{-1}$, driven by strong SOC and broken TRS yang2023tuning . In Fe3GeTe2 monolayers, an intrinsic AHC of $\approx$350 $\Omega^{-1}\text{cm}^{-1}$ has been reported, with further enhancement achieved through external tuning methods such as ionic-liquid gating guo2023modulating . Similarly, studies on MnBi2Te4 have observed intrinsic AHC values ranging from 250 to 300 $\Omega^{-1}\text{cm}^{-1}$, attributed to contributions from Berry curvature and canted antiferromagnetic ordering. To the best of our knowledge, there are no previous reports on AHC in 2D kagome materials. Our predicted intrinsic AHC value of $\approx$257 $\Omega^{-1}\text{cm}^{-1}$ falls within the range of values reported for other 2D systems. Moreover, it is consistent with the values observed in bulk kagome materials, which range from $\sim$380 $\Omega^{-1}\text{cm}^{-1}$ for bulk LiMn6Sn6 chen2021large to as high as 500 $\Omega^{-1}\text{cm}^{-1}$ for Fe3Sn2 belbase2023large .
In Figure 3, we investigate the impact of an externally applied perpendicular electric field on the topological properties of 2D ScV6Sn6. The edge states (Figures 3(a)–(b)), associated with the Weyl point located near the Fermi energy, and the Berry curvature distribution at zero field are characterized by topological band crossings along the MMKK high-symmetry path in the Brillouin zone. Notably, the Berry curvature hotspots (Figure 3(b)) coincide with Weyl point pairs both at and slightly below the Fermi level within the valence band. Upon applying an external electric field, the electronic structure undergoes systematic evolution. While low-intensity fields cause minimal perturbation, preserving the Weyl-like crossings and the overall topological character, higher field strengths induce significant band rearrangements. In particular, the chirality and location of Weyl points shift progressively, and a critical field strength of approximately 0.4 eV/Å marks the onset of a topological phase transition. At this threshold, spin–orbit coupling-driven bulk gaps emerge, and the previously gapless edge states begin to vanish, as shown in Figure 3(f). The disappearance of topologically protected edge modes, along with a change in the Chern number from 1 to 0 (Figure 3(e)), unambiguously confirms a transition from a nontrivial Weyl semimetal to a trivial metallic phase. Beyond this critical field, enhanced electronic coherence manifests through the formation of fully interconnected Fermi pockets (Figure 3(d)), indicative of delocalized charge carriers in the trivial regime. Similar electric-field-induced topological transitions have been observed in other 2D materials, including MnBi2Te4, MoS2, and WSe2, where applied fields serve as effective tuning parameters to drive phase transitions between trivial and topological states Qian_2014 ; Sawahata_2019 ; you2021electric . Although the system remains metallic across the entire range of applied electric field strengths, this field-induced evolution highlights a controllable pathway for tuning topological phases in 2D materials. Achieving a perpendicular electric field on the order of 0.4 eV/Å may be experimentally challenging via conventional gating methods; however, comparable field magnitudes can be realized through ionic-liquid gating in field-effect transistor architectures. Alternatively, substrate engineering employing polar or ferroelectric materials offers a promising route to modulate the local electrostatic potential and thereby manipulate the underlying band topology weintrub2022generating ; lv2019reconfigurable .
@@
In our simulations, we observed fluctuations in the Chern number at smaller electric fields, likely caused by modifications in the Berry curvature, driven by electric-field-induced alterations in the band structure. Although the precise mechanism behind these fluctuations remains unclear, similar behavior has been reported in MnBi2Te4, where the Chern number can be toggled by applying an electric field, with transitions between discrete values such as 1 and 2 at field strengths of around 1 to 2 eV/Å Cai_2022 . Additionally, thin films near the topological critical point have exhibited fluctuations in the Chern number between 0, 1, and -1 as the electric field changes, eventually stabilizing at zero beyond approximately 2.5 eV/Å lin2015quantum . Rasouli et al. also observed similar electric-field-induced changes in the Chern number for 2D MnO2, with stabilization at zero occurring beyond a certain threshold.
@@
In contrast, the application of small magnetic fields (B-fields) does not alter the topology of the bands. The Chern number remains unchanged regardless of the direction or magnitude of the applied magnetic field. This stability is due to the robustness of the topologically protected edge states, preserved by the system’s geometry and time-reversal symmetry (TRS). As moderate magnetic fields are insufficient to break this symmetry, the topological phase remains protected, similar to the role of spin-orbit coupling (SOC), which is also unable to break the symmetry moore2010birth . Moreover, the weak coupling between the magnetic field and the electronic states in this material results in minimal effects on the band structure and topological properties. The stability of the Fermi surface under both in-plane and out-of-plane magnetic fields up to 2.1 T (not shown) further supports this robustness. The Fermi surface retains its 2D character, characterized by large hexagonal V-$d$ states and smaller circular Sc-$d$ states, as illustrated in Figures 3(b)-(c). Importantly, the shape of the Fermi surface remains consistent across different magnetic field strengths, indicating that the electronic structure is preserved under external magnetic fields. These findings are consistent with studies on other magnetic topological materials. For example, in MnBi2Te4, although magnetic fields induce transitions between antiferromagnetic (AFM) and ferromagnetic (FM) states, the Fermi surface remains largely unaffected, and no topological transition is observed tan2022mnbi2te4 . Similarly, in Cr-doped Bi2Se3 and 2D materials such as CrI3 and Cr2Ge2Te2, minimal topological changes occur under magnetic fields despite notable effects on magnetic properties nano13192655 ; liu2024magnetoresistance .
@@ IV Conclusions
We have investigated the 2D kagome metal ScV6Sn6 and demonstrated that it exhibits topologically protected edge states and a Chern number of $|\textit{C}|=1$ at the Fermi energy, confirming the presence of a Weyl semimetallic phase. This positions ScV6Sn6 among the few vanadium-based kagome compounds with nontrivial topology and broken time-reversal symmetry. Our analysis of edge states originating from gapped Weyl points near the Fermi energy reveals that the material retains metallic behavior at the surface despite being gapped in the bulk, confirming its topological nature. We further show that while magnetic fields do not affect its topology, an electric field of 0.4 eV/Å induces a transition from a topological semimetal to a trivial metallic phase, with the edge states vanishing. This electrically induced transition, consistent with observations in other 2D materials, underscores the tunability of ScV6Sn6’s topological properties. Additionally, our calculations reveal a large anomalous Hall effect of $\sim$257 $\Omega^{-1}$cm-1, driven by intrinsic Berry curvature and strong spin-orbit coupling, aligning with other 2D ferromagnetic Dirac-like materials. These findings highlight ScV6Sn6 as a pivotal material in 2D kagome systems, with strong potential for modulating topological properties in $d$-orbital-dominated structures. Its robust ferromagnetic ordering, tunable topological phases, and spin-orbit coupling make it an ideal candidate for advanced electronic and spintronic applications. The ability to control topological transitions through electric fields not only enhances our understanding of topological quantum materials but also paves the way for innovations in quantum computing and material design.
@@ V Acknowledgments
This research is supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences under Award DOE-SC0024099. SKD was supported by the start up grant of Bitan Roy from Lehigh University. Computational resources were provided by Lehigh University Research Computing Infrastructure. We also acknowledge useful discussions with Benjamin Wieder.